# New black hole mergers in the LIGO–Virgo O3 data from a gravitational wave search including higher-order harmonics

### v1

@@ II.3.3 Ranking statistic
Even after performing the above steps, there are still some noise triggers which survive. We use the ranking statistic as a final tool to downweight such triggers. Beyond upweighting triggers which have good coherence among multiple detectors, an important job of the ranking statistic is to also downweight triggers based on the probability with which similar triggers occur in the background. We simulate 2000 instances of background runs using the method of timeslides (where we shift one of the detectors by unphysical amounts which are more than the light crossing time between the detectors).
In the -only case, for a trigger with SNR associated with a template , this probability under the noise hypothesis is . In the case of Gaussian noise, follows a (i.e., ). However, we empirically measure by constructing a histogram of the values of background triggers associated with the template . In our previous papers, we labelled this term as the “rank function” (), see section J of Ref. Venumadhav et al. (2019).
@@
In a similar way for the HM case, if the noise were Gaussian, separately follow distributions (note that we orthogonalize templates of different harmonics), and their sum thus follows a distribution with 6 d.o.f. This motivated us to make the approximation and again measure this probability distribution empirically by constructing a 1D histogram of values of background triggers associated with the template . We plan to explore beyond the current assumption in a future study.
@@
Note that in the previous IAS studies, is assumed to be the same for all templates in a subbank. However, especially for high masses, this term can have different values even for different templates within a subbank, see e.g., Abbott et al. (2016b); Nitz et al. (2017); Wadekar et al. (2023b). In order to optimize our analysis, we follow the technique in Ref. Wadekar et al. (2023b), where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template). We then separately construct the histograms of background triggers for each template group to estimate the noise probabilities of triggers associated to templates within that group.
In a future paper, we plan to quantify the improvement in volume-time sensitivity due to the background reduction procedures outlined in this section, by injecting simulated gravitational-wave signals in the data and comparing the fraction of injections recovered by our pipeline.
@@
1 | GW190605_025957 | 13,0 | ||||||||||||
2 | GW190806_033721 | 8,1 | ||||||||||||
3 | GW190524_134109 | 10,0 | ||||||||||||
4 | GW191113_103541 | 9,0 | ||||||||||||
5 | GW190615_030234 | 10,0 | ||||||||||||
6 | GW190604_103812 | 12,0 | ||||||||||||
7 | GW191228_195619 | 14,0 | ||||||||||||
8 | GW200304_172806 | 8,1 | ||||||||||||
9 | GW190530_030659 | 4,2 | ||||||||||||
10 | GW190511_163209 | 8,0 | ||||||||||||
11 | GW200210_100022 | 9,0 | ||||||||||||
12 | GW200301_211019 | 1,2 | ||||||||||||
13 | GW190911_195101 | 12,0 | ||||||||||||
14 | GW190530_133833 | 12,0 |
@@
GW200109_195634 | 10,0 | ||||||||||||
GW190818_232544 | 8,1 | ||||||||||||
GW190906_054335 | 7,1 |
GWTC-3 events | Bank | IFAR (yr) | ||||||||
@@
GW190828_063405 | BBH_4 | |||||||||
GW190828_065509 | Veto (BBH_1) | 50.14 | 56.48 | |||||||
GW190909_114149111GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. | BBH_7 | |||||||||
GW190915_235702 | BBH_4 | |||||||||
GW190916_200658 | BBH_8 | |||||||||
GW190917_114630 | Below | |||||||||
@@
GW191109_010717 | Veto (BBH_8) | 82.74 | 164.37 |
GWTC-3 events | Bank | IFAR (yr) | ||||||||
@@
GW200322_091133 | BBH_6 | |||||||||
4-OGC events | Bank | IFAR (yr) | ||||||||
@@
GW200318_191337 | BBH_7 | 63.0 |
II.4 Astrophysical prior
where refer to detector-frame (redshifted) masses of the binary. We chose the location of the mass ratio cut for BBHs to restrict to the calibration regime of the IMRPhenomXAS approximant Pratten et al. (2020), while the cut on the individual spins is due to the relativistic Kerr limit.
We picked a simple prior which broadly covers the parameter space (picking a broad prior can help with reweighting the results to a different prior in the future). Our prior on the mass ratio is a uniform one on for ; the effective spin parameter is flat within (with ); the total mass is distributed in a power-law fashion, . Note that we use a prior on the detector-frame masses, and we do not assume a separate prior over redshift .
We drew samples according to this prior (each sample corresponding to the physical parameter set ), and assigned a weight to each sample based on its observable volume. As a proxy for the observable distance, we simply calculated the SNR for the sample at a fixed distance by simulating its waveform using IMRPhenomXHM and using the same reference PSD which was used to make the template banks (weight SNR). Just for reference, for inspiral dominated waveforms, the observable volume roughly scales as Fishbach and Holz (2017). Using these samples, we then utilized the kernel density estimation method to construct probabilities for templates in our template banks (see appendix C of O22 for details on the density estimation method). Fig. 7 in the Appendix shows the cumulative probabilities for our banks, each of which is obtained by adding the weights of all the samples falling within the corresponding bank.
II.4.1 Comparison with astrophysical prior used in our -only banks
It is worth briefly comparing the astrophysical distributions we assumed here to those we assumed in our previous -only O3 searches. We split our banks roughly according to the detector-frame chirp mass (the bank edges were uniform in log ) and further split each bank into subbanks based on the amplitude profiles of normalized waveforms Roulet et al. (2019). We assumed that the total prior astrophysical probability within each of our banks was identical, and that the probability distribution within each bank was flat in the detector-frame component masses and also in the effective spin (without taking the observable distance into account; we take this effect into account in the current search). Note that we also slightly increased our search space in this paper: previously, our banks were limited to and (see figure 2 of O22), while we now use .
III Results
III.1 New candidate events
Using our new search pipeline on the Hanford–Livingston data from the third observing run (O3), we find 14 candidate events with which have not been previously reported in the literature. We list the properties of these new candidates in Table 1. In all tables in this paper, , columns correspond to incoherent SNR obtained from in the particular detector (note that we have orthogonalized the templates of different harmonics, so that they can be added in quadrature). The incoherent values may not correspond to those possible for a physical trigger. Thus, in some cases we see that can even exceed the best-fit parameter estimation (PE) values . Further discussion of the PE runs are given below in Section III.1.1, and we report 90% CL values of some of the binary parameters in Table 1. We show the SNR contribution of individual modes separately in Table 5 in the Appendix.
We also mention the best bank_id, subbank_id for each event; the bank_id roughly increases with mass (see Fig. 7). Note that the high-mass banks in our case have much fewer templates than the low-mass banks (see Table I of Wadekar et al. (2023a)), which leads to considerable heterogeneity in the number of background triggers in the banks. Hence, the values needed for events to be above threshold can be different for different banks.
where in the above formula corresponds to our final ranking statistic value for the candidate events (see Ref. Wadekar et al. (2023b) for the full expression), and corresponds to number density values from a 1D histogram of the ranking statistic (which we separately calculate for the signal and background cases). Note that as the templates in our banks have astrophysical probabilities associated with them (see Section II.4), and we use these values in calculating . Hence, our IFAR and values are also dependent on our assumption of astrophysical distributions of binary parameters. We expect this dependence to be more significant for the ‘overall’ IFAR in comparison to the ‘per bank’ IFAR values quoted in Table 1. We therefore separately provide the ‘per bank’ values in our table, so one can reweight them in a rough way for different astrophysical distributions than the ones assumed here.
@@ III.1.1 Parameter estimation runs
We perform parameter estimation (PE) runs using the package222https://github.com/jroulet/cogwheel Roulet et al. (2022) with the sampler Lange (2023). We use priors that are uniform in detector-frame constituent masses and comoving volume-time () throughout this paper. Our fiducial prior for the BH spins is flat in effective spin , but we also compare results with those obtained using an isotropic prior on individual spins (which is used for PE in the GWTC-3 and 4-OGC catalogs). Other extrinsic parameters have the standard geometric priors used in GWTC-3 Abbott et al. (2021). We compute likelihoods using the relative binning/heterodyning method described in Ref. Zackay et al. (2018) as implemented in cogwheel. The relation between redshift and luminosity distance is based on the Planck15 cosmology Ade et al. (2016). Our fiducial parameter estimation runs are based on the IMRPhenomXODE waveform approximant Yu et al. (2023) and include the effect of both HM and precession. Table 1 reports marginalized 1D posterior values for a relevant subset of parameters.
@@
Other than the fiducial runs, we also perform two separate PE runs without precession, i.e., restricting to binaries with aligned spins: one run with a -only waveform approximant and the other with an approximant that includes HM as well. This helps us calculate the relative evidence (Bayes factor) due to including the effects of HM and HM + precession separately for the new candidate events; we show the logarithm of the Bayes factors in Table 1. Four of the new candidate events (the ones numbered ) have Bayes factors larger than unity for HM, and the detection of these candidates likely benefited the most from our search containing HM. Most of the remaining events have Bayes factors for HM that are lower than unity, and we judge that the detection of these additional candidates was a result of just an overall improvement of our search’s sensitivity in the high mass/redshift region due to e.g., improved template coverage and background estimation.
@@
We show a few select 2D posteriors of our new candidate events in colored contours in Figs. 1, 2 and 7. We also overplot the contours from events in GWTC-3, 4-OGC and the IAS -only O3 catalogs in light gray. Our search only covers events with Hanford–Livingston coincidence triggers, but we include all binary BH events that are part of the LVK catalogs through O1–O3 in these figures. Note that the posteriors are bi-modal for some events, which show up as two contours in the plot.
@@
We add dashed lines to mark the approximate limits of sensitivity in the top panel of Fig. 1. The dashed line in the top-right of the panel marks the region where the binary starts moving out of band (we take this to happen when the detector frame mass equals , which is the limit of our template bank). The dashed line in the top-left marks the contour of constant detector-frame luminosity for an inspiral-dominated equal-mass binary; it roughly delineates the parameter limits past which binaries are too faint to be observable. It is also worth mentioning that the size of the contours increases for higher masses as the number of cycles in band starts to decrease. In the inspiral phase, the chirp mass is well-measured, but this no longer continues to be the case for high-mass events where the SNR is dominated by the merger and ringdown phase and the total mass of the binary starts to be measured better.
We show the marginalized 1D posteriors for our new candidate events under the aforementioned two different spin priors in Fig. 3. Figs. 10 and 11 in the Appendix show full corner plots from different PE runs. Some of the events in Fig. 3 (e.g., the ones numbered ) have posteriors that show significant support for large, positive effective spin under the uniform in prior. The isotropic spin prior recovers a lower likelihood for these events, since the prior pulls the sampler away from the high solution.
@@
GW191228_195619 has the largest evidence for HM () and also for the combination of HM+precession (). We show a corner plot corresponding to this event in Fig. 4 for our three PE runs, and we indeed see significant support for both HM and precession. Note that for the (green) and + HM (orange) cases, the distribution of the primary component’s in-plane spin component is that of its prior conditioned on the other parameters. For these cases, we set the in-plane spin components to zero when evaluating likelihoods. Within 90% CL, we find to be . Note also that the total SNR (110) recovered by our non-precessing Hanford+Livingston search is significantly smaller than obtained from a PE run which includes both the effects of orbital precession and data from the Virgo detector. Therefore, we conjecture that the significance of this event could be improved in future searches that include the effects of precession and/or Virgo data.
@@
We also compare our results against PE runs using the IMRPhenomXPHM approximant Pratten et al. (2021) (under the two different MSA Chatziioannou et al. (2017) and NNLO Hannam et al. (2014) prescriptions for evolving the spins) in Fig. 9 in the Appendix; we find good agreement in most of the cases. It would be interesting to compare the PE for these events with NR surrogate models (e.g. Islam et al. (2023)) in the future.
We will discuss interesting astrophysical properties of our new candidate events in Section IV below.
@@ III.2 Comparison to the IAS (2, 2)-only catalogs
In Table 2, we report the candidate events with from our pipeline which have already been reported in the previous -only IAS searches in the O3 data Olsen et al. (2022); Mehta et al. (2023). The full comparison catalog including lower candidates is given in Table 6. Overall, we recover 6 events with .
We caution against directly comparing the sensitivities of the -only and + HM pipelines from Table 6 as there are multiple differences: () including HM in the search, () using a different astrophysical prior (see Section II.4) () additional glitch cleaning at high masses (see Section II.3). It is worth mentioning though that we recover the events which were reported in previous IAS catalogs with , except GW190704_104834 (this event was in BBH-0 in O22 and we checked that the difference in the astrophysical priors between the pipelines was large for that bank). We collect a larger number of triggers in the HM search as compared to a -only search, so we raised the collection threshold to mitigate the increased memory requirements. This caused us to miss low-SNR events that were reported by -only pipelines with lower collection thresholds in the previous literature. We further discuss this point in Section III.3.2 below.
One notable event in Table 2 is GW190711_030756. The IFAR of this event in our HM search is 38.7 yr as opposed to yr in the -only search O22 (after accounting for the trials factor in O22 due to the number of banks searched). We show a corner plot for this event in Fig. 5. Indeed, GW190711_030756 also shows a significant evidence for higher modes and our search recovers an additional 10 points in SNR compared to the -only search. See Fig. 3 and section III A.2 of O22 for a detailed discussion of the properties of this event.

### v2

@@ II.3.4 Ranking statistic
Even after performing the above noise-removal steps, there are still some noise triggers which survive. Our ranking statistic is the final tool we can use to downweight such triggers. Beyond upweighting triggers which have good coherence among multiple detectors, an important job of the ranking statistic is to also downweight triggers based on the probability with which similar triggers occur in the background. We simulate 2000 instances of background runs using the method of timeslides (where we shift one of the detectors by unphysical amounts which are more than the light crossing time between the detectors).
@@
In the -only case, for a trigger with SNR associated with a template , this probability under the noise hypothesis is . In the case of Gaussian noise, follows a (i.e., ). However, we empirically measure by constructing a histogram of the values of background triggers associated with the template . In a similar way for the HM case, if the noise were Gaussian, separately follow distributions (note that we orthogonalize templates of different harmonics), and their sum thus follows a distribution with 6 d.o.f. This motivated us to make the approximation and again measure this probability distribution empirically by constructing a 1D histogram of values of background triggers associated with the template (see W24).
@@
Note that in the previous IAS studies, is assumed to be the same for all templates in a subbank. However, especially for high masses, this term can have different values even for different templates within a subbank, see e.g., Abbott et al. (2016b); Nitz et al. (2017); Wadekar et al. (2024). In order to optimize our analysis, we follow the technique in W24, where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template). We then separately construct the histograms of background triggers for each template group to estimate the noise probabilities of triggers associated to templates within that group.
@@
| 1 | GW200104_184028 | 8,1 | | | | | ||||||||
| 2 | GW190425_133124 | 14,0 | | | | | ||||||||
| 3 | GW200319_022728 | 7,1 | | | | | ||||||||
| 4 | GW190605_025957 | 14,0 | | | | | ||||||||
| 5 | GW200114_020818 | 8,1 | | | | | ||||||||
| 6 | GW190524_134109 | 9,0 | | | | | ||||||||
| 7 | GW191213_204950 | 1,2 | | | | | ||||||||
| 8 | GW190530_030659 | 3,2 | | | | | ||||||||
| 9 | GW200210_035448 | 9,0 | | | | | ||||||||
| 10 | GW190908_161547 | 7,1 | | | | | ||||||||
| 11 | GW190710_150742 | 10,0 | | | | | ||||||||
II.4 Astrophysical prior
where refer to detector-frame (redshifted) masses of the binary. We chose the mass ratio cut for BBHs to restrict to the calibration regime of the IMRPhenomXAS approximant to numerical relativity simulations Pratten et al. (2020).
We picked a simple prior which broadly covers the parameter space (picking a broad prior can help with reweighting the results to a different prior in the future Roulet et al. (2020)). We use a uniform prior on the mass ratio for ; the effective spin parameter is flat within (with ); the total mass is distributed in a power-law fashion, . Note that we use a prior on the detector-frame masses, and we do not assume a separate prior over redshift .
We drew samples according to this prior (each sample corresponding to the physical parameter set ), and assigned a weight to each sample based on its observable volume. We improve upon the astrophysical prior used in O22 by incorporating the fact that the observable changes across the parameter space. For example, for inspiral dominated waveforms, the observable volume roughly scales as Fishbach and Holz (2017). As a proxy for the observable distance, we simply calculated the SNR for the sample at a fixed distance by simulating its waveform using IMRPhenomXHM under a reference PSD (weightsample SNR3).
Using these samples, we then utilized normalizing flows to predict the effective astrophysical prior density at the location of our templates. We also use normalizing flows to generate physical and samples corresponding to each template (which are required by our marginalization integral). We leave further discussion of our use of normalizing flows to an upcoming paper Wadekar et al. (b). Fig. 7 in the Appendix shows the cumulative probabilities for our banks, each of which is obtained by adding the weights of all the samples falling within the corresponding bank.
III Results
III.1 New candidate events
Using our new search pipeline on the Hanford–Livingston data from the third observing run (O3), we find 11 candidate events with which have not been previously reported in the literature. We list the properties of these new candidates in Table 1. In all tables in this paper, , columns correspond to incoherent SNR obtained from in the particular detector (note that we have orthogonalized the templates of different harmonics, so that they can be added in quadrature). The incoherent values may not correspond to those possible for a physical trigger. Thus, in some cases we see that can even exceed the best-fit parameter estimation (PE) values . Further discussion of the PE runs are given below in Section III.1.1, and we report 90% CL values of some of the binary parameters in Table 1. We show the SNR contribution of individual modes separately in Table 5 in the Appendix.
We also mention the best bank_id, subbank_id for each event; the bank_id roughly increases with mass (see Fig. 7). Note that the high-mass banks in our case have much fewer templates than the low-mass banks (see Table I of Wadekar et al. (2023)), which leads to considerable heterogeneity in the number of background triggers in the banks. Hence, the values needed for events to be above threshold can be different for different banks.
where in the above formula corresponds to our final ranking statistic value for the candidate events (see Eq. 16 of Ref. W24 for the full expression), and corresponds to number density values from a 1D histogram of the ranking statistic (which we separately calculate for the signal and background cases). Note that as the templates in our banks have astrophysical probabilities associated with them (see Section II.4), and we use these values in calculating . Hence, our IFAR and values are also dependent on our assumption of astrophysical distributions of binary parameters. We expect this dependence to be more significant for the ‘overall’ IFAR in comparison to the ‘per bank’ IFAR values quoted in Table 1. We therefore separately provide the ‘per bank’ values in our table, so one could reweight them in a rough way for different astrophysical distributions than the ones assumed here.
@@ III.1 New candidate events
| GW191117_023843 | 10,0 | | | | | | | | |||||
| GWTC-3 events | Bank | GWTC-3 | IFAR (yr) | |||||||
@@
| GW190828_063405 | BBH_3 | |||||||||
| GW190828_065509 | BBH_1 | |||||||||
| GW190909_114149222GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. | BBH_5 | |||||||||
| GW190915_235702 | BBH_3 | |||||||||
| GW190916_200658 | BBH_6 | |||||||||
| GW190917_114630 | BBH_0 | |||||||||
@@
| GW191109_010717 | BBH_7 | |||||||||
| GWTC-3 events | Bank | GWTC-3 | IFAR (yr) | |||||||
@@
| GW200322_091133 | BBH_1 | |||||||||
| 4-OGC events | Bank | 4-OGC | IFAR (yr) | |||||||
@@
| GW200318_191337 | BBH_7 | 63.0 | ||||||||
@@ III.1.1 Parameter estimation runs
We perform parameter estimation (PE) runs using the package333https://github.com/jroulet/cogwheel Roulet et al. (2022) with the sampler Lange (2023). We use priors that are uniform in detector-frame constituent masses and comoving volume-time () throughout this paper. Our fiducial prior for the BH spins is flat in effective spin , but we also compare results with those obtained using an isotropic prior on individual spins (which is used for PE in the GWTC-3 and 4-OGC catalogs). Other extrinsic parameters have the standard geometric priors used in GWTC-3 Abbott et al. (2021). We compute likelihoods using the relative binning/heterodyning method described in Ref. Zackay et al. (2018); Roulet et al. (2024) as implemented in cogwheel. The relation between redshift and luminosity distance is based on the Planck15 cosmology Ade et al. (2016). Our fiducial parameter estimation runs are based on the IMRPhenomXODE waveform approximant Yu et al. (2023) and include the effect of both HM and precession. Table 1 reports marginalized 1D posterior values for a relevant subset of parameters.
@@
Other than the fiducial runs, we also perform two separate PE runs without precession, i.e., restricting to binaries with aligned spins: one run with a -only waveform approximant and the other with an approximant that includes HM as well. This helps us calculate the relative evidence (Bayes factor) due to including the effects of HM and HM + precession separately for the new candidate events; we show the logarithm of the Bayes factors in Table 1. Two of the new candidate events have Bayes factors larger than unity for HM, and the detection of these candidates likely benefited the most from our search containing HM.
@@
We show a few select 2D posteriors of our new candidate events in colored contours in Figs. 1, 2 and 6. We also overplot the contours from events in GWTC-3, 4-OGC and the IAS -only O3 catalogs in light gray. Our search only covers events with Hanford–Livingston coincidence triggers, but we include all binary BH events that are part of the LVK catalogs through O1–O3 in these figures.
@@
We add dashed lines to mark the approximate limits of sensitivity in the top panel of Fig. 1. The dashed line in the top-right of the panel marks the region where the binary starts moving out of band (we roughly take this to happen when the detector frame mass equals , which is the limit of our template bank). The dashed line in the top-left marks the contour of constant detector-frame luminosity for an inspiral-dominated equal-mass binary; it roughly delineates the parameter limits past which binaries are too faint to be observable. It is also worth mentioning that the size of the contours increases for higher masses as the number of cycles in band starts to decrease. In the inspiral phase, the chirp mass is well-measured, but this no longer continues to be the case for high-mass events where the SNR is dominated by the merger and ringdown phase and the total mass of the binary starts to be measured better.
@@
We show the marginalized 1D posteriors for our new candidate events under the aforementioned two different spin priors in Fig. 3. Figs. 9 and 10 in the Appendix show full corner plots from different PE runs. We additionally highlight the posteriors for two most interesting new candidate events (GW200114_020818, GW190425_133124) in Fig. 5. GW200114_020818 has the largest evidence for HM (). We show three PE runs with and without HM and precessing effects. Note that for the non-precessing (green and orange) cases, the distribution of the in-plane spin component is same as that of the prior conditioned on the other parameters (we set contribution of in-plane spin components to be zero inside the likelihood for these cases). Note also that the total SNR2 (180) recovered by our non-precessing Hanford+Livingston search is significantly smaller than obtained from a PE run which includes both the effects of orbital precession and data from the Virgo detector. Similarly, SNR77 and for GW190425_133124. Therefore, we predict that the significance of these event could be improved in future searches that include the effects of precession and/or Virgo data.
@@
We also compare our results against PE runs using the IMRPhenomXPHM approximant Pratten et al. (2021) (using the default MSA Chatziioannou et al. (2017) prescription for evolving the spins) in Fig. 8 in the Appendix; we find good agreement in most of the cases. It would be interesting to compare the PE for these events with NR surrogate models (e.g. Islam et al. (2023)) in the future. We will discuss interesting astrophysical properties of our new candidate events in Section IV below.
@@ III.2 Comparison to the IAS (2, 2)-only catalogs
In Table 2, we report the candidate events with from our pipeline which have already been reported in the previous -only IAS searches in the O3 data Olsen et al. (2022); Mehta et al. (2023). The full comparison catalog including lower candidates is given in Table 6. Overall, we recover 5 events with .
We caution against directly comparing the sensitivities of the -only and + HM pipelines from Table 6 as there are multiple differences: () including HM in the search, () using a different astrophysical prior (see Section II.4) () additional glitch mitigation at high masses (see Section II.3). It is worth mentioning though that we recover the events which were reported in previous IAS catalogs with , except GW190704_104834 (this event was in BBH-0 in O22 and we checked that the difference in the astrophysical priors between the pipelines was large for that bank). We collect a larger number of triggers in the HM search as compared to a -only search, so we raised the collection threshold to mitigate the increased memory requirements (see Section II.2.1). This caused us to miss low-SNR events that were reported by -only pipelines with lower collection thresholds in the previous literature.
One notable event in Table 2 is GW190711_030756. The IFAR of this event in our HM search is 11 yr as opposed to yr in the -only search O22 (after accounting for the trials factor in O22 due to the number of banks searched). We show a corner plot for this event in Fig. 4. Indeed, GW190711_030756 also shows a significant evidence for higher modes and our search recovers an additional 10 points in SNR2 compared to the -only search. See Fig. 3 and section III A.2 of O22 for a detailed discussion of the properties of this event.